OpenStudy (littlenugget):

Find the domain of the function algebraically and support your answer graphically

OpenStudy (littlenugget):

\[\frac{ 3x-1 }{ (x+3)(x-1) }\]

OpenStudy (anonymous):

What do you mean

OpenStudy (anonymous):

To find the domain, you would have to look at the denominator and see what X cannot equal, or else it will be undefined. Set x+3=0 Set x-1=0 Whatever equals zero, X cannot equal. These are excluded from your domain.

OpenStudy (anonymous):

Ok

OpenStudy (debbieg):

I think what is meant by the "support your answer graphically" part is to show that the graph of the function has vertical asymptotes at those values of x that are not in the domain.

OpenStudy (littlenugget):

So my teach like taught us how to do it graphically with the calculator, but not algebraically. So basically you're saying that x=-3 and x=1 can not be in the domain? @dimensionx

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Because you cannot have zero in the denominator.

OpenStudy (littlenugget):

so then the domain would be\[(-\infty,-4) \upsilon (2,\infty)\]

OpenStudy (littlenugget):

@dimensionx

OpenStudy (debbieg):

How did you get that? @dimensionx told you to solve: x+3=0 x-1=0 What do you get?

OpenStudy (littlenugget):

x=-3 and x=1

OpenStudy (debbieg):

OK, so if you leave those two values of x out of the domain, then what do you get? (I'm not seeing how you came up with \((-\infty,-4) \cup (2,\infty)\)??)

OpenStudy (debbieg):

First off, you are leaving out all the values between -4 and 2.... but also, -4 and 2 aren't the values that are to be left out of the domain.

OpenStudy (littlenugget):

so it would be \[(-\infty,-3)\upsilon(1,\infty)\]

OpenStudy (littlenugget):

am i close? lol

OpenStudy (debbieg):

OK, that's better - closer, yes! :) But you are leaving out everything between -3 and 1. EVERYTHING except for -3 and 1 are in the domain.

OpenStudy (littlenugget):

so is it like \[(-\infty,-3)\upsilon(-3,1)\upsilon(1,\infty)\]

OpenStudy (debbieg):

YES, exactly like that. :) good job!

OpenStudy (littlenugget):

oh yeahhhh!!!!!!! :DDDDD so it's pretty easy lol thank you so much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! @DebbieG can i like double fan you? lol

OpenStudy (debbieg):

lol you're welcome... happy to help :)

OpenStudy (anonymous):

{R}-(-3) {R}-1